Paula rolls a fair number cube with sides labeled through 6. what is the probability of Paula rolling that is 3 or greater


1
Expert's answer
2022-01-25T04:44:43-0500

Four numbers are 3 or greater. 3,4,5, and 6.


"P(X\geq3)=\dfrac{4}{6}=\dfrac{2}{3}"

The probability of Paula rolling that is 3 or greater is "\dfrac{2}{3}."